# Geodesic Slice Sampling on the Sphere

### v4

@@ 4.3 Curved Distribution on the Sphere
For our initial sampling tests for the curved vMF target, we consider $\mathbb{S}^{2}$ and choose $\kappa=500$. Again, we run all four MCMC samplers. Rejection-based geoSSS and HMC take roughly the same amount of time, with HMC taking slightly longer. However, both are almost twice as slow as shrinkage-based geoSSS. RWMH is by far the fastest method (refer to Appendix I for run times).
Figure 8: First panel from left shows the curved vMF target density on the $2$-sphere with concentration parameter $\kappa=500$. Remaining panels show the last 5000 samples from the four different MCMC samplers targeting this density.
Figure 8 shows the target density and samples of the four MCMC algorithms generated in the last 5000 iterations. The underlying target density shown here is evaluated by uniformly distributing $3\times 10^{4}$ points on the $2$-sphere using the method of Saff1997. Visual inspection shows that the slice samplers and also HMC explore the target much more rapidly than RWMH, which is also reflected in the high autocorrelation (see Fig. 9A). To verify this observation, we measure the discrepancy between the discretized target and the histogram generated by the sampling procedure using KL divergence (see Fig. 9B). Moreover, we monitor the distance between successive samples (see Fig. 9C). Across all three evaluation we consider, RWMH is clearly disfavored. On the other hand, geoSSS and HMC show a similar performance on this low-dimensional target.
Figure 9: (A) ACF of the first dimension plotted for the curved vMF target on $2$-sphere based on $10^{5}$ samples. (B) KL divergence between a discretized version of the curved vMF target on 2-sphere and a histogram compiled from the MCMC samples. (C) Geodesic distance (log-scale) between successive approximate samples for the curved von Mises-Fisher distribution on $2$-sphere (17).
@@ 5 Summary
All authors are grateful for the support of the DFG within project 432680300 – SFB 1456 subprojects A05 and B02. We also thank Philip Schär for valuable comments and discussions about the topic. M. Habeck and S. Kodgirwar gratefully acknowledge funding by the Carl Zeiss Foundation within the program “CZS Stiftungsprofessuren” and by the German Research Foundation (DFG) within grant HA 5918/4-1.
@@
We have implemented our geoSSS algorithms, including the spherical RWMH and HMC methods (detailed in Appendices F and G), in our geosss package available at https://github.com/microscopic-image-analysis/geosss. You can find the scripts used to generate the numerical illustrations in the scripts directory, and the protein example data in the data directory. Also the precomputed results are archived on Zenodo at https://doi.org/10.5281/zenodo.15610462.
@@ Appendix D Proof of Lemma 11
In this section we prove the integral estimate of Lemma 11. A major part of deriving it consists of handling the measure resulting from exploring the sphere along the great circle passing through a fixed $x\in\mathbb{S}^{d-1}$. At first sight, one could think that this measure is $\sigma_{d-1}$. Keeping in mind though that, interpreting $x$ as the “north pole”, the great circle lie much “denser” at the poles than at the equator, it becomes rather clear that $\sigma_{d-1}$ can only be a lower estimate for the measure obtained by exploring $\mathbb{S}^{d-1}$ via the great circle through $x$.
@@
and the absolute value of their Jacobians
@@
$\int_{0}^{\uppi}\int_{{\mathbb{S}^{d-2}_{x}}}\mathbbm{1}_{A}\big{(}\gamma_{(x,v)}(\theta)\big{)}\,\mu_{x}({\rm d}v)\,{\rm d}\theta=\int_{\mathbb{S}^{d-1}}\mathbbm{1}_{A}\left(\widetilde{\varphi}_{x}(y)\right)g(y)\,\sigma_{d-1}({\rm d}y).$
@@
$\int_{0}^{\uppi}\int_{{\mathbb{S}^{d-2}_{x}}}\mathbbm{1}_{A}\big{(}\gamma_{(x,v)}(\theta)\big{)}\,\mu_{x}({\rm d}v)\,{\rm d}\theta\geq\int_{\mathbb{S}^{d-1}}\mathbbm{1}_{A}(\widetilde{\varphi}_{x}(y))\,\sigma_{d-1}({\rm d}y)=\sigma_{d-1}(A).$
@@
Appendix E Proof of Proposition 15.
@@
$\begin{split}Q_{{x,}{v,}{t}}(\cdot)=\mathbb{P}\big{(}\text{shrink}(x,v,t)\in\cdot\big{)}=\mathbb{P}\big{(}\overline{\text{shrink}}(0,L(x,v,t))\in\cdot\big{)}=\bar{Q}_{L(x,v,t)}(0,\cdot),\end{split}$ (23)
@@
Appendix F Random-walk Metropolis-Hastings on the Sphere.
The random-walk Metropolis-Hastings (RWMH) algorithm uses an isotropic Gaussian proposal kernel. As suggested by Lie et al. (2023), for current state $x\in\mathbb{S}^{d-1}$ we choose an auxiliary point $\sqrt{r}x$ in the ambient space $\mathbb{R}^{d}$ by generating a radius $\sqrt{r}$, with $r$ being a realization of $R\sim{\rm Gamma}(d/2,1/2)\,.$ Then, given $\sqrt{r}x$ we sample a realization $y$ from $Y\sim{\rm Normal}\left(\sqrt{r}x,\varepsilon^{2}\mathrm{Id}_{d}\right),$ where $\varepsilon$ is the step-size of the random walk. Since this $y$ does not yet lie on the sphere, we radially project and propose $y/\|y\|$, which finally is accepted or rejected using the usual acceptance ratio. In Algortihm 7 we provide the corresponding pseudocode.
@@
4: Draw $U\sim\mathcal{U}(0,1)$, call the result $u$.
@@
Appendix G Hamiltonian Monte Carlo on the Sphere.
Since RWMH suffers from diffusive behavior, we also test spherical Hamiltonian Monte Carlo (HMC) suggested by Lan et al. (2014) as an alternative MCMC approach. In spherical HMC, the sample space is first augmented by momenta or velocities $v\in\mathbb{R}^{d}$ that are in the the tangent space $E_{x}$ to $\mathbb{S}^{d-1}$ at the current sample $x$. The momenta follow a standard Normal distribution, and the Markov chain is generated in the space $\{(x,v)\mid x\in\mathbb{S}^{d-1},v\in E_{x}\}$. Diffusive behavior is suppressed by the special type of proposal step that solves Hamilton’s equations of motion by using a leapfrog integrator (Neal, 2011). During leapfrog integration, the gradient of the log probability $\nabla\log p(x)$ guides the Markov chain, thereby reaching nearby modes in much shorter time than RWMH. The spherical HMC algorithm is detailed in Algorithm 8. In contrast to standard HMC, spherical HMC involves a rotation of the positions and momenta during leapfrog integration where the rotation matrix is a Givens rotation $G(v/\|v\|,x,\theta)$ in the plane spanned by the momenta $v$ and the positions $x$ (see lines 8–10 in Algorithm 8 and (4)). In addition to the step-size parameter $\varepsilon>0$, we also need to choose the number of leapfrog steps $T\in\mathbb{N}$. In our experiments, we always set $T=10$.
@@ Appendix H Step-size Tuning for RWMH and HMC
10: Draw $U\sim\mathcal{U}(0,1)$, call the result $u$.
@@ Appendix I Run Times of Numerical Illustrations
We summarize the run times in seconds (s) for all the numerical experiments tested on Intel i7-10510U CPU in Table A1. In addition to the tests reported in Section 4, we provide a Jupyter notebook for testing the geodesic slice samplers on Bingham targets in various dimensions.111github.com/microscopic-image-analysis/geosss/scripts/Bingham.ipynb We also report run times for these targets. For brevity we name these experiments “Bingham-d10” for the Bingham distribution with $d=10$, $\kappa_{d}=30$ and $N=10^{5}$; “Bingham-d50” for the Bingham distribution with $d=50$, $\kappa_{d}=300$ and $N=10^{5}$; “vMF-d10” for the mixture of vMF distribution with $d=10$, $\kappa=100$, $K=5$ and $N=10^{6}$; “curved-vMF-d3” for the curve vMF distribution with $d=3$, $\kappa=500$ and $N=10^{6}$; “curved-vMF-d5” for curve vMF with $d=5$, $\kappa=800$ and $N=10^{6}$.
Numerical MCMC Methods
Illustrations geoSSS (reject) geoSSS (shrink) RWMH HMC
@@
vMF-d10 2601.6 s 812.9 s 274.4 s 1898.9 s
curved-vMF-d3 17106.9 s 5583.2 s 1841.0 s 12292.9 s
curved-vMF-d5 42483.7 s 5509.0 s 1355.7 s 8659.0 s
Table A1: Comparison of run times for each MCMC method across different numerical illustrations.

### v5

@@ 4.3 Curved Distribution on the Sphere
For our initial sampling tests for the curved vMF target, we consider $\mathbb{S}^{2}$ and choose $\kappa=500$. Again, we run all four MCMC samplers. Rejection-based geoSSS and HMC [^3] More details on the implementation of the required gradient can be found in the Python library https://github.com/microscopic-image-analysis/geosss and in the following note geosss/scripts/gradient-curve.pdf take roughly the same amount of time, with HMC taking slightly longer. However, both are almost twice as slow as shrinkage-based geoSSS. RWMH is by far the fastest method (refer to Appendix I for run times).
Figure 8: First panel from the left shows the curved vMF target density on the $2$-sphere with concentration parameter $\kappa=500$. Remaining panels show the last 5000 samples from the four different MCMC samplers targeting this density.
Figure 8 shows the target density and samples of the four MCMC algorithms generated in the last 5000 iterations. The underlying target density shown here is evaluated by uniformly distributing $3\times 10^{4}$ points on the $2$-sphere using the method of Saff and Kuijlaars (1997). Visual inspection shows that the slice samplers and also HMC explore the target much more rapidly than RWMH, which is also reflected in the high autocorrelation (see Fig. 9A). To verify this observation, we measure the discrepancy between the discretized target and the histogram generated by the sampling procedure using KL divergence (see Fig. 9B). Moreover, we monitor the distance between successive samples (see Fig. 9C). Across all three evaluation criteria we consider, RWMH is clearly disfavored. On the other hand, geoSSS and HMC show a similar performance on this low-dimensional target.
Figure 9: (A) ACF of the first dimension plotted for the curved vMF target on $2$-sphere based on $10^{5}$ samples. (B) KL divergence between a discretized version of the curved vMF target on 2-sphere and a histogram compiled from the MCMC samples. (C) Geodesic distance (log-scale) between successive approximate samples for the curved von Mises-Fisher distribution (17) on 22-sphere.
@@ 5 Summary
All authors are grateful for the support of the DFG within project 432680300 – SFB 1456 subprojects A05 and B02. We also thank Philip Schär for valuable comments and discussions about the topic. M. Habeck and S. Kodgirwar gratefully acknowledge funding by the Carl Zeiss Stiftung within the program “CZS Stiftungsprofessuren” and by the German Research Foundation (DFG) within grant HA 5918/4-1. M. Hasenpflug and D. Rudolf were supported by the German Research Foundation (DFG grant number 522337282).
@@
We have implemented our geoSSS algorithms, including the spherical RWMH and HMC methods (detailed in Appendices F and G), in our geosss package available at https://github.com/microscopic-image-analysis/geosss. You can find the scripts used to generate the numerical illustrations in the scripts directory, and the protein example data in the data directory. Also the precomputed results are archived on Science Data Bank https://doi.org/10.57760/sciencedb.30181.
@@ Appendix D Proof of Lemma 11
In this section we prove the integral estimate of Lemma 11. A major part of deriving it consists of handling the measure resulting from exploring the sphere along the great circle passing through a fixed $x\in\mathbb{S}^{d-1}$. At first sight, one could think that this measure is $\sigma_{d-1}$. Keeping in mind though that, interpreting $x$ as the “north pole”, the great circles lie much “denser” at the poles than at the equator, it becomes rather clear that $\sigma_{d-1}$ can only be a lower estimate for the measure obtained by exploring $\mathbb{S}^{d-1}$ via the great circles through $x$.
@@
and the absolute value of the determinant of their Jacobians
@@
$\int_{0}^{\uppi}\int_{{\mathbb{S}^{d-2}_{x}}}\mathbbm{1}_{A}\big(\gamma_{(x,v)}(\theta)\big)\,\mu_{x}({\rm d}v)\,{\rm d}\theta=\int_{\mathbb{S}^{d-1}}\mathbbm{1}_{A}\big(\widetilde{\varphi}_{x}(y)\big)g(y)\,\sigma_{d-1}({\rm d}y).$
@@
$\int_{0}^{\uppi}\int_{{\mathbb{S}^{d-2}_{x}}}\mathbbm{1}_{A}\big(\gamma_{(x,v)}(\theta)\big)\,\mu_{x}({\rm d}v)\,{\rm d}\theta\geq\int_{\mathbb{S}^{d-1}}\mathbbm{1}_{A}\big(\widetilde{\varphi}_{x}(y)\big)\,\sigma_{d-1}({\rm d}y)=\sigma_{d-1}(A).$
@@
Appendix E Proof of Proposition 15
@@
$\begin{split}Q_{{x,}{v,}{t}}(\cdot)&=\mathbb{P}\big(\text{shrink}(x,v,t)\in\cdot\mod 2\uppi\big)=\mathbb{P}\big(\overline{\text{shrink}}(0,L(x,v,t))\in\cdot\big)\\ &=\bar{Q}_{L(x,v,t)}(0,\cdot),\end{split}$ (23)
@@
Appendix F Random-walk Metropolis-Hastings on the Sphere
The random-walk Metropolis-Hastings (RWMH) algorithm uses an isotropic Gaussian proposal kernel. As suggested by Lie et al. (2023), for current state $x\in\mathbb{S}^{d-1}$ we choose an auxiliary point $\sqrt{r}x$ in the ambient space $\mathbb{R}^{d}$ by generating a radius $\sqrt{r}$, with $r$ being a realization of $R\sim{\rm Gamma}(d/2,1/2)\,.$ Then, given $\sqrt{r}x$ we sample a realization $y$ from $Y\sim{\rm Normal}\left(\sqrt{r}x,\varepsilon^{2}\mathrm{Id}_{d}\right),$ where $\varepsilon$ is the step-size of the random walk. Since this $y$ does not yet lie on the sphere, we radially project and propose $y/\|y\|$, which finally is accepted or rejected using the usual acceptance ratio. In Algorithm 7 we provide the corresponding pseudocode. Note that we also introduce mixture-MH, which can be interpreted as a variant that combines RWMH with independent Metropolis. For details we refer to Algorithm 8, which is specifically employed for testing the mixture of vMF distributions.
@@ Appendix F Random-walk Metropolis-Hastings on the Sphere
4: Draw $U\sim\mathcal{U}_{(0,1)}$, call the result $u$.
@@
Algorithm 8 Mixture of RWMH and Independence Sampler on $\mathbb{S}^{d-1}$ with mixture hyperparameter $\alpha\in[0,1]$ and step-size $\varepsilon>0$.
input: current state $x\in\mathbb{S}^{d-1}$
output: next state $x^{\prime}$
1: Draw $U_{1}\sim\mathcal{U}_{(0,1)}$, call the result $u_{1}$.
2:if $u_{1}<\alpha$ then
3:  // Use random walk proposal
4:  Draw $R\sim{\rm Gamma}\left(d/2,1/2\right)$, call the result $r$.
5:  Draw $Y\sim{\rm Normal}\left(\sqrt{r}\,x,\varepsilon^{2}\mathrm{Id}_{d}\right)$, call the result $y$.
6:  Set $z=y/\|y\|$.
7:else
8:  // Use independent proposal
9:  Draw $Y\sim{\rm Normal}\left(0,\mathrm{Id}_{d}\right)$, call the result $y$.
10:  Set $z=y/\|y\|$.
11:end if
12: Draw $U_{2}\sim\mathcal{U}_{(0,1)}$, call the result $u_{2}$.
13:if $u_{2}\leq\min\left\{1,p(z)/p(x)\right\}$ then
14:  $x^{\prime}=z$
15:else
16:  $x^{\prime}=x$
17:end if
Appendix G Hamiltonian Monte Carlo on the Sphere
Since RWMH suffers from diffusive behavior, we also test spherical Hamiltonian Monte Carlo (HMC) suggested by Lan et al. (2014) as an alternative MCMC approach. In spherical HMC, the sample space is first augmented by momenta or velocities $v\in\mathbb{R}^{d}$ that are in the the tangent space $E_{x}$ to $\mathbb{S}^{d-1}$ at the current sample $x$. The momenta follow a standard Normal distribution, and the Markov chain is generated in the space $\{(x,v)\mid x\in\mathbb{S}^{d-1},v\in E_{x}\}$. Diffusive behavior is suppressed by the special type of proposal step that solves Hamilton’s equations of motion by using a leapfrog integrator (Neal, 2011). During leapfrog integration, the gradient of the log probability $\nabla\log p(x)$ guides the Markov chain, thereby reaching nearby modes in much shorter time than RWMH. The spherical HMC algorithm is detailed in Algorithm 9. In contrast to standard HMC, spherical HMC involves a rotation of the positions and momenta during leapfrog integration where the rotation matrix is a Givens rotation $G(v/\|v\|,x,\theta)$ in the plane spanned by the momentum $v$ and the position $x$ (see lines 8–10 in Algorithm 9 and Equation 4). In addition to the step-size parameter $\varepsilon>0$, we also need to choose the number of leapfrog steps $T\in\mathbb{N}$. In our experiments, we always set $T=10$.
@@ Appendix H Step-size Tuning for RWMH and HMC
10: Draw $U\sim\mathcal{U}_{(0,1)}$, call the result $u$.
@@ Appendix I Run Times of Numerical Illustrations
We summarize the run times in seconds (s) for all the numerical experiments tested on Intel i7-10510U CPU in Table A1. In addition to the tests reported in Section 4, we provide a Jupyter notebook for testing the geodesic slice samplers on Bingham targets in various dimensions.444github.com/microscopic-image-analysis/geosss/scripts/Bingham.ipynb We also report run times for these targets. For brevity we name these experiments “Bingham-d10” for the Bingham distribution with $d=10$, $\kappa_{d}=30$ and $N=10^{5}$; “Bingham-d50” for the Bingham distribution with $d=50$, $\kappa_{d}=300$ and $N=10^{5}$; “vMF-d10” for the mixture of vMF distribution with $d=10$, $\kappa=100$, $K=5$ and $N=10^{6}$; “curved-vMF-d3” for the curve vMF distribution with $d=3$, $\kappa=500$ and $N=10^{6}$; “curved-vMF-d5” for curve vMF with $d=5$, $\kappa=800$ and $N=10^{6}$.
Experiments geoSSS (reject) geoSSS (shrink) RWMH HMC
@@
vMF-d10* 2601.6 s 812.9 s 274.4 s 1898.9 s
curve-vMF-d3 17106.9 s 5583.2 s 1841.0 s 12292.9 s
curve-vMF-d5 42483.7 s 5509.0 s 1355.7 s 8659.0 s
    The extensively tuned mixture-MH sampler employed only for this test took 577.0 s.
Table A1: Comparison of run times for each MCMC method across different numerical experiments.